OpenStudy (anonymous):

Gravel is being dumped from a conveyor belt at a rate of 20 cubic feet per minute. It forms a pile in the shape of a right circular cone whose base diameter and height are always equal. How fast is the height of the pile increasing when the pile is 21 feet high? Recall that the volume of a right circular cone with height h and radius of the base r is given by V=1/3πr^(2)h

OpenStudy (anonymous):

@Chlorophyll ok so first start by deriving the function (2 pi h r)/3

OpenStudy (anonymous):

wow that's tough

OpenStudy (anonymous):

then if im not missing anything if you plugged in (2 pi (21) (20))/3 but i feel as if thats way to easy to be right

OpenStudy (anonymous):

volume is 1/3 *pi *r2*h

OpenStudy (anonymous):

and ya it is tough thats why im on here

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Since the "diameter and height are always equal " -> 2r = h => r = h/2 Now plug r into the volume v. We utilize the relationship to obtain only 1 variable h: V= 1/3π r² h

OpenStudy (anonymous):

@NickR Questions?

OpenStudy (anonymous):

V=1/12 *pi *h3

OpenStudy (anonymous):

differentiate with respect to time

OpenStudy (anonymous):

u know dv/dt=20 cubic/min

OpenStudy (anonymous):

plug in to the equation dV/dt=1/4* pi *h2*dh/dt

OpenStudy (anonymous):

h=21 dv/dt=20

OpenStudy (anonymous):

um so V= 1/3π (h/2)² h

OpenStudy (anonymous):

dh/dt=0.057

OpenStudy (anonymous):

Yup, convert the equation into 1indepent variable prior to taking derivative!

OpenStudy (anonymous):

plz give me medal

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i will tell this to my mom that i got a medal

OpenStudy (anonymous):

(pi h^2)/4 (pi 21^2)/4

OpenStudy (anonymous):

@Fazeelayaz you're truly deserved it :)

OpenStudy (anonymous):

@NickR Following @Fazeelayaz post staring from : V = ( π/12 ) h³

OpenStudy (anonymous):

*starting from:

OpenStudy (anonymous):

20=1/4* pi *21^2*.057

OpenStudy (anonymous):

=> h' = 80/ 21² π

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

Yes, h' = .057 ft/min

OpenStudy (anonymous):

oh ok didnt know pi was in denominator

OpenStudy (anonymous):

Feel free to ask questions :)

OpenStudy (anonymous):

ok sorry to keep bothering you but got one more question if you could look at it and this time i actually know how to do it.

OpenStudy (anonymous):

Okie, open new post :)